OpenStudy (anonymous):

MEDAL Solve the system by the method of your choice. Identify systems with no solution and systems with infinitely many solutions, using set notation to express their solution sets. x + y = -8 x - y = 14

OpenStudy (anonymous):

if you choose to add the two equations, what will happen ?

OpenStudy (anonymous):

x + y = -8 --- (1) x - y = 14 --- (2) (1) - (2): x + y - x + y = -8 - 14 2y = -22 y = -11 Substitute the value of y into (1): x - 11 = -8 x = 3 Thus, x = 3, y = -11.

OpenStudy (anonymous):

Didn't she say by "using set notation to express their solution sets"?

OpenStudy (anonymous):

thank you !